Question with last attempt is displayed for your review only The time it takes me to wash the dishes is uniformly distributed between 7 minutes and 13 minutes. What is the probability that washing dishes tonight will take me between 9 and 12 minutes? Give your answer accurate to two decimal places. Incorrect Box 1: Enter your answer as an integer or decimal number. Examples: 3, -4, 5.5172 Enter DNE for Does Not Exist, oo for Infinity
Answer #1

Solution :

Given that,

a = 7

b = 13

P(c < x < d) = (d - c) / (b - a)

P(9 < x < 12) = (12 - 9) / (13 - 7)=0.50

probability=0.50